OpenStudy (anonymous):

A car enters an interstate highway 20 mi north of a city. The car travels north at an average speed of 58 mph. Which equation models the car’s distance d from the city after traveling for h hours?

OpenStudy (anonymous):

what are the options?

OpenStudy (anonymous):

these are my choices... A. 20 d = h + 58 B. d= 58 h + 20 C. d= 20 h + 58 D. 58 = 20 h + d

OpenStudy (anonymous):

How many miles does the car travel?

OpenStudy (anonymous):

Read the question @PeytonPerez372

OpenStudy (anonymous):

so the answer would be A.???? @RosieF

OpenStudy (anonymous):

YES!!!!!! great job :)

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

NP!! :)